Imagine:

Solve Correctly 23x+2=117

umm:

Given\[23x+2=117\]Subtract 2 from both sides\[23x+2−2=117−2\]\[23x=115\]Divide both sides by 23\[\frac{ 23x }{ 23 }=\frac{ 115 }{ 23 }\]\[x=5\]Therefore your answer is\[x=5\]

Imagine:

wow that is way better explaining.